OpenStudy (anonymous):

last year there was 23 eighth graders now there are 217 find percent increase

OpenStudy (anonymous):

sorry it is 203 people not 23

OpenStudy (anonymous):

217-203=14 14/203=x/10 cross and time 14*100=203x x= 1400/203 x=6.896 or 0.07%

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

thank you sooooooo much

OpenStudy (anonymous):

u r we com